OpenStudy (anonymous):

what formula do you use to find the value of x in a pair of similar figures. The figures are 1.5ft & x 3ft & 6ft

OpenStudy (anonymous):

for similar fig. 1.5/x =3/6 x=2*1.5=3(by cross multiplication)

OpenStudy (amistre64):

Lets call the numbers L1=1.5 and L2=x; R1=3 and R2=6 You can set this up a few ways. L1/L2 = R1/R2 L1/R1 = L2/R2 L2/L1 = R2/R1 and so forth. I tend to use the most basic: L1/L2 = R1/R2 1.5/x = 3/6 1.5 = 3x/6 1.5/3 = x/6 1.5 (6/3) = x 1.5 (2) = x 3=x